I'm looking for the name of a theorem, and the completion of the theorem. I've been searching the internet and my algebra books, but can't locate the theorem. The question marks is what I need to fill in. If you know the theorem, I can look it up to complete the statement.

The ? Theorem states if p and q are real numbers and if p^2 = q^2, then ?
The ? Theorem states if p and q are real numbers and if p * q = 0, then ?

Thanks!

Answer by Edwin McCravy(20056)

I'm looking for the name of a theorem, and the completion of the theorem. I've been searching the internet and my algebra books, but can't locate the theorem. The question marks is what I need to fill in. If you know the theorem, I can look it up to complete the statement.
The ? Theorem states if p and q are real numbers and if p^2 = q^2, then ?

The ? Theorem states if p and q are real numbers and if p * q = 0, then ?
Thanks!

I can answer only three of the four:

The ????? Theorem states if p and q are real numbers and if p^2 = q^2, then 
|p| = |q| 


The zero-factor Theorem states if p and q are real numbers and if p * q = 0,
then either p=0 or q=0 or both.

Sorry, but I do not know of any name for the first. My guess would be
the "absolute value theorem", but that's only a guess. 

Edwin